# GraphBench: Next-generation graph learning benchmarking

### v1

@@ Results
In Table˜19 we provide results for each weather variable at three different pressure levels, aligning with the results obtained in WeatherBench2 (rasp2024weatherbench), which allows for an easily accessible comparison of weather forecasting models with traditional weather forecasting baselines. For comparative models like GraphCast and Persistence, only selected results are available. Thus, we compare against those in particular. However, we also account for cumulative errors across all pressure levels in the underlying data. Throughout our evaluation, we provide a 12-hour forecast prediction across all variables. Our baselines establish a transparent and reproducible lower bound for medium-range weather forecasting. While they do not reach the performance of Persistence or specialized systems like GraphCast, this is expected given their simpler architecture, shorter training time, and deliberately straightforward setup. Several factors contribute to this outcome. Unlike specialized systems, our models were trained for substantially fewer iterations, with fewer computing resources and without extensive hyperparameter optimization. We deliberately prioritized transparency and reproducibility over complexity in the architectural design. Our training setup is intentionally simple and does not incorporate domain-specific refinements, such as positional edge features, which are known to improve forecast skill. Moreover, medium-range weather forecasting is a particularly demanding task that often benefits from specialized evaluation protocols, whereas our assessment relies on a basic, uniform metric. Importantly, the purpose of this baseline is not to achieve state-of-the-art performance. Its main contribution is to provide a transparent, reproducible lower bound that serves as a precise reference point against which future graph learning methods can be compared.
@@ 4 Format, licensing, and long-term access
All GraphBench datasets are distributed as PyG objects with graph data, and raw data is also available in HDF5 for long-term compatibility. Data is released under open licenses, i.e., Apache 2.0 for most datasets (including WeatherBench2), MIT for SAT instances, and GPL for AClib. Regarding social networks, we report that the original BlueSky data are hosted at https://zenodo.org/records/11082879 under the Creative Commons Attribution 4.0 International License. The owners of such data may update these records to comply to users’ “Right to Erasure”; and we are in the process of implementing a system to update our derived dataset consistently. GraphBench is continuously updated with new datasets, and data is hosted on institutional servers and public repositories for persistent access.
@@ 5 Automated hyperparameter optimization
GraphBench also integrates automated hyperparameter optimization (HPO), which aims to easily tune new models, improve performance, and enhance reproducibility. The hyperparameter optimization in GraphBench is based on the SMAC3 (LinEtAl22) package, which utilizes multi-fidelity scheduling. This approach evaluates many configurations on lower budgets and eliminates those with poor performance. As graph learning approaches are expensive to train, this approach enables the saving of costly resources. Additionally, SMAC3 utilizes a surrogate model to propose new configurations, thereby learning from previous experience and further reducing the computational load. Due to limited computational resources, we were unable to run HPO on all datasets and domains. To showcase the effectiveness of the automated hyperparameter optimization included in GraphBench, we tuned the GIN model on the small SAT dataset with VG. We set the total optimization budget to 150 evaluations, varying fidelity over training gradient steps between 1 000 and 100 000. The used configuration space is shown in Table˜20, and the results are presented in Table˜21. We observe a $7{\ziffer@DotOri}3\%$ reduction in RMSE between the manually tuned and automatically tuned versions, demonstrating that automated HPO in GraphBench can improve model performance.
@@ 6 Conclusion and Outlook
Appendix A Appendix
Appendix B Dataset Details
@@
B.1 Social Networks: BlueSky
@@
Appendix C Experimental setup
@@
where $\sigma$ denotes the GeLU nonlinearity and $\mathbold{W}_{1}\ziffer@CommaOri\mathbold{W}_{2}$ are weight matrices. Further, $\phi$ is the selected baseline, and NNN denotes a two-layer feed-forward neural network with GELU non-linearity (Hendrycks2016Gelu). For the GT baseline, we employed biased attention as described by Spe+2025. Furthermore, we incorporate RWSE (Dwivedi2022RWSE) and LPE (Mueller2024Aligning) as node positional encodings for our MPNN and GT baselines.
@@
Using the aforementioned hyperparameter tuning process, we evaluate all baseline models across three random seeds per task and report their mean performance along with their standard deviations. For out-of-distribution generalization tasks, we utilize pretrained models from these baseline evaluations.
@@ C.1 Hyperparameter selection
C.2 Baseline Architectures
In the following, we provide implementation details on the baselines used for the datasets in GraphBench. For dataset-specific design choices, we provide detailed information in Section˜C.3.
@@ C.2 Baseline Architectures
As described in Appendix˜C we use an encoder-processor-decoder baseline across tasks. In this case, we consider a GT as the processor architecture following the implementation outlined in Spe+2025. For most tasks, we use node-level tokenization, treating each graph node as a single token input to the GT. However, for edge-level tasks, we use the transformation outlined for algorithmic reasoning tasks, allowing edge-level tokens to be used without changes to the processor architecture. Additionally, absolute PEs, such as RWSE or LPE, are added to the node embeddings before the first GT layer. Then, the GT layer computes full multi-head scaled-dot-product attention, adding an attention bias $\mathbold{B}$ to the unnormalized attention matrix and applying softmax. Let $\mathbold{Q}\ziffer@CommaOri\mathbold{K}\ziffer@CommaOri\mathbold{V}\in\mathbb{R}^{L\times d}$ and $\mathbold{B}\in\mathbb{R}^{L\times L}$ with $L$ denoting the number of tokens and $d$ the embedding dimension. Then, attention and a GT layer take the form
@@
where $\mathbold{W}_{Q}\ziffer@CommaOri\mathbold{W}_{K}\ziffer@CommaOri\mathbold{W}_{V}\in\mathbb{R}^{d\times d}$ are learnable weight matrices and $\mathsf{MLP}$ denotes a two layer-MLP. In practice, we compute attention over multiple heads, allowing for different attention biases to be added to the attention matrix. With $\phi$ given by the multihead attention computation, a processor layer is provided by Equation˜3. In the processor, multiple layers are stacked, allowing the pipeline shown in Appendix˜C.
@@
Throughout this work, we use a GINE-based graph neural network baseline as the processor in the framework outlined in Appendix˜C. Following, HuStrategiesPretraining the GINE layer updates the node representations $h_{v}^{(t)}$ at iteration $t$ as follows:
@@
We apply the GINE processor layer in a similar way to the GT baseline design outlined previously by replacing $\phi$ with a GINE layer where $\mathsf{N}$ is given by a dropout layer. First, the node embeddings, optionally with added PEs, are passed to the GINE layer, where layer normalization is applied. The output of the GINE message-passing layer is then forwarded to a two-layer MLP. We use the same residual connection as seen in the GIN implementation from Spe+2025. We then stack multiple layers together to provide the processor part of our baseline architecture. Unless otherwise specified, we use mean pooling for graph-level tasks at the end of the processor step.
C.3 Architecture choices for datasets
@@
where $\sigma$ is set to ReLU, dropout is applied before its application and, finally, no normalization layers are interleaved.

### v2

@@ Results
In Table˜19 we present results for each weather variable at three pressure levels, aligning with the results obtained in WeatherBench2 (rasp2024weatherbench), enabling an easily accessible comparison of weather forecasting models with traditional weather forecasting baselines. For comparative models like GraphCast and Persistence, only selected results are available. Thus, we compare against those in particular. However, we also account for cumulative errors across all pressure levels in the underlying data. Throughout our evaluation, we provide a 12-hour forecast prediction across all variables. Our baselines establish a transparent and reproducible lower bound for medium-range weather forecasting. While they do not reach the performance of Persistence or specialized systems like GraphCast, this is expected given their simpler architecture, shorter training time, and deliberately straightforward setup. Several factors contribute to this outcome. Unlike specialized systems, our models were trained for substantially fewer iterations, with fewer computing resources and without extensive hyperparameter optimization. We deliberately prioritized transparency and reproducibility over complexity in the architectural design. Our training setup is intentionally simple and does not incorporate domain-specific refinements, such as positional edge features, which are known to improve forecast skill. Moreover, medium-range weather forecasting is a particularly demanding task that often benefits from specialized evaluation protocols, whereas our assessment relies on a basic, uniform metric. Importantly, the purpose of this baseline is not to achieve state-of-the-art performance. Its main contribution is to provide a transparent, reproducible lower bound that serves as a precise reference point against which future graph learning methods can be compared.
@@ 4 Format, licensing, and long-term access
All GraphBench datasets are distributed as PyG objects with graph data, and raw data is also available in HDF5 for long-term compatibility. Data is released under open licenses, i.e., Apache 2.0 for most datasets (including WeatherBench2), MIT for SAT instances, and GPL for AClib. Regarding social networks, we report that the original BlueSky data are hosted at https://zenodo.org/records/11082879 under the Creative Commons Attribution 4.0 International License. The owners of such data may update these records to comply with users’ “Right to Erasure”, and we are in the process of implementing a system to update our dataset consistently. GraphBench is continuously updated with new datasets, and data is hosted on institutional servers and public repositories for persistent access.
@@ 5 Automated hyperparameter optimization
GraphBench also integrates automated hyperparameter optimization (HPO), which aims to easily tune new models, improve performance, and enhance reproducibility. The hyperparameter optimization in GraphBench is based on the SMAC3 (LinEtAl22) package, which utilizes multi-fidelity scheduling. This approach evaluates many configurations on lower budgets and eliminates those with poor performance. As graph learning approaches are expensive to train, this approach enables the saving of costly resources. Additionally, SMAC3 uses a surrogate model to propose new configurations, thereby leveraging prior experience and further reducing computational load. Due to limited computational resources, we were unable to run HPO on all datasets and domains. To showcase the effectiveness of the automated hyperparameter optimization included in GraphBench, we tuned the GIN model on the small SAT dataset with VG. We set the total optimization budget to 150 evaluations, varying fidelity over training gradient steps between 1 000 and 100 000. The used configuration space is shown in Table˜20, and the results are presented in Table˜21. We observe a $7{\ziffer@DotOri}3\%$ reduction in RMSE between the manually tuned and automatically tuned versions, demonstrating that automated HPO in GraphBench can improve model performance.
@@ 6 Conclusion and outlook
Appendix A Dataset details
@@
A.1 Social networks: BlueSky
@@
Appendix B Experimental setup
@@
where $\sigma$ denotes the GeLU nonlinearity and $\mathbold{W}_{1}\ziffer@CommaOri\mathbold{W}_{2}$ are weight matrices. Further, $\phi$ is the selected baseline, and NNN denotes a two-layer feed-forward neural network with GELU non-linearity [Hendrycks2016Gelu]. For the GT baseline, we employed biased attention as described by Spe+2025. Furthermore, we incorporate RWSE [Dwivedi2022RWSE] and LPE [Mueller2024Aligning] as node positional encodings for our MPNN and GT baselines.
@@
Using the aforementioned hyperparameter tuning process, we evaluate all baseline models across three random seeds per task and report their mean performance and standard deviations. For out-of-distribution generalization tasks, we utilize pretrained models from these baseline evaluations.
@@ B.1 Hyperparameter selection
B.2 Baseline architectures
In the following, we provide implementation details on the baselines used for the datasets in GraphBench. For dataset-specific design choices, we provide detailed information in Section˜B.3.
@@ B.2 Baseline architectures
As described in Appendix˜B we use an encoder-processor-decoder baseline across tasks. In this case, we consider a GT as the processor architecture following the implementation outlined in Spe+2025. For most tasks, we use node-level tokenization, treating each graph node as a single token input to the GT. However, for edge-level tasks, we use the transformation outlined for algorithmic reasoning tasks, allowing edge-level tokens to be used without changes to the processor architecture. Additionally, absolute PEs, such as RWSE or LPE, are added to the node embeddings before the first GT layer. Then, the GT layer computes full multi-head scaled-dot-product attention, adding an attention bias $\mathbold{B}$ to the unnormalized attention matrix and applying softmax. Let $\mathbold{Q}\ziffer@CommaOri\mathbold{K}\ziffer@CommaOri\mathbold{V}\in\mathbb{R}^{L\times d}$ and $\mathbold{B}\in\mathbb{R}^{L\times L}$ with $L$ denoting the number of tokens and $d$ the embedding dimension. Then, attention and a GT layer take the form
@@
where $\mathbold{W}_{Q}\ziffer@CommaOri\mathbold{W}_{K}\ziffer@CommaOri\mathbold{W}_{V}\in\mathbb{R}^{d\times d}$ are learnable weight matrices and $\mathsf{MLP}$ denotes a two layer-MLP. In practice, we compute attention over multiple heads, allowing for different attention biases to be added to the attention matrix. With $\phi$ given by the multihead attention computation, a processor layer is provided by Equation˜3. In the processor, multiple layers are stacked, allowing the pipeline shown in Appendix˜B.
@@
Throughout this work, we use a GINE-based graph neural network baseline as the processor in the framework outlined in Appendix˜B. Following, HuStrategiesPretraining the GINE layer updates the node representations $h_{v}^{(t)}$ at iteration $t$ as follows:
@@
We apply the GINE processor layer in a similar way to the GT baseline design outlined previously, by replacing $\phi$ with a GINE layer with $\mathsf{N}$ determined by a dropout layer. First, the node embeddings, optionally with added PEs, are passed to the GINE layer, where layer normalization is applied. The output of the GINE message-passing layer is then forwarded to a two-layer MLP. We use the same residual connection as seen in the GIN implementation from Spe+2025. We then stack multiple layers together to provide the processor part of our baseline architecture. Unless otherwise specified, we use mean pooling for graph-level tasks at the end of the processor step.
B.3 Architecture choices for datasets
@@
where $\sigma$ is set to ReLU, dropout is applied before its application, and, finally, no normalization layers are interleaved.